14. + 3.58/7.18 points Previous Answers TanFin 11 5.3.048. My Notes Five years ago, Diane secured a bank loan of $340,000 toHelp
Answer #1

(c) Outstanding Mortgage Balance at the end of Year 5 = $ 316384.81

New Tenure = 30 years or (30 x 12) = 360 months

Interest Rate = 2.5 % compounded monthly, Applicable Monthly Rate = 2.5/12 = 0.2083 %

Let the new required monthly repayments be $ m

Therefore, 316384.81 = m x (1/0.002083) x [1-{1/(1.002083)^(360)}]

316384.81 = m x 253.087

m = 316384.81 / 253.087 = $ 1250.1

(d) Original Mortgage Repayments = $ 2038.47 and New Mortgage Repayments = $ 1250.1

Monthly Savings = 2038.47 - 1250.1 = $ 788.37